OpenStudy (anonymous):

Find the 10th term of the sequence whose nth term is given by the formula tn = (-1)n+1/n. A. -1/10 B. -1/7 C. 1/7 D. 1/10

OpenStudy (anonymous):

you should rewrite this. The present notation is too vague

OpenStudy (anonymous):

Truh, none of the answers is right the way you wrote it.

OpenStudy (anonymous):

Thats how my school wrote it...

OpenStudy (anonymous):

sorry Find the 10th term of the sequence whose nth term is given by the formula tn = (-1)^n+1/n. A. -1/10 B. -1/7 C. 1/7 D. 1/10

OpenStudy (anonymous):

something is still off

OpenStudy (anonymous):

Yes, something still missing.

OpenStudy (anonymous):

That is exactly how it looks on my homework

OpenStudy (anonymous):

are we supposed to be solving for t? or is that t(n) where n is a function of t

OpenStudy (anonymous):

You should recheck your homework again. There is an error somewhere (either answer bank or notation)

OpenStudy (anonymous):

Its horrible quality but here is a picture of my homework

OpenStudy (anonymous):

there it is (n+1) is the exponent

OpenStudy (anonymous):

\[t _{10}=(-1)^{10+1}/10=\frac{ -1 }{ 10 }\] the trick to see here is quickly tell if the exponent will yield a positive 1 or negative 1. -1 raised to an even power is positive while -1 raised to an odd power will be negative. so we can know (-1)^102938449942002 will be = 1 since the exponent is even

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

look again @Partycool

OpenStudy (anonymous):

sorry A got mixed up